a) calculate the normality of the solution , if 50 g of caustic soda is dissolved in water to prepare 2000 ml of solution

b) 86 g of potassium carbonate is dissolved in 200 g of water .the density of the solutin is 1.5 g/cc . find the molarity, molality and normality .

c) calculate the volume of the tank in cm3 to store 45 kg of ethane at a temperature of 50 c and pressure of 50 atm . consider ethane to be an ideal gas
1
Expert's answer
2020-05-04T01:53:10-0400

Pa) m(NaOH) = 50g

V(solution) = 2000 mL

N(solution) -?

eq(NaOH) = 1

N = m/(Meq*V) = 50g/(40g/mol* 2L) = 0.625 N

b) m(K2CO3) = 86 g

V(H2O) = 200 g

ρ = 1.5 g/cc = 1500 g/L

M - ? m - ? N -?

ρ = m/V

V = m/ρ = (86 + 200)/1500 = 0.19 L

M(solution) = n(K2CO3)/ V(solution) = 3.28 M

m(solution) = n(K2CO3)/ m(solvent) = 3.12 mol/kg

Meq(K2CO3) = 69 g/mol*eq

N(solution) = m(K2CO3)/(Meq(K2CO3)*V) = 86 g/ (69 g/mol*eq * 0.19 L) = 6.56 N

c) m(C2H6) = 45 kg = 45000 g

t = 500C + 273 = 323 K

P = 50 atm = 50.5*105Pa

V - ? (cm3)

P*V = m*R*T/M

V = m*R*T/P*M = 45000 g*8.31 J/mol*K *323 K / 50.5*105Pa *30g/mol = 0.80 m3 =

= 8*105cm3